OpenStudy (anonymous):

let f(x)=-2x+4 and g(x)=-6x-7. find f(x)-g(x). Please show work, I don't get it. The other problem I don't get is f(x)=3x-7 and g(x)=-2x-6. Find (f o g)(4). I don't understand any of this, finished the other questions.

OpenStudy (mathmale):

If you want to find f(x)-g(x), write f(x) first and g(x) underneath it, and then subtract, column by column.

OpenStudy (mathmale):

let f(x)=-2x+4 and g(x)=-6x-7. find f(x)-g(x). f(x)=-2x+4 - g(x)=-6x-7 -------------\ Combine each column.

OpenStudy (anonymous):

Thanks, but what about the second problem. And I finished the other one. =-8x-3 correct? @mathmale